OpenStudy (anonymous):

I would greatly appreciate any help! Thanks so much (2x(1/3))/1-(1/3)^2

OpenStudy (anonymous):

confused a bit. looks like \[\frac{2x \times \frac{1}{3}}{1-(\frac{1}{3})^2}\] but that doesn't really make sense.

OpenStudy (anonymous):

perhaps the 'x' is a times sign?

OpenStudy (anonymous):

lol oh yes, I'm sorry. The x was a multiplication sign.

OpenStudy (anonymous):

ah so it is \[\frac{2\times \frac{1}{3}}{1-(\frac{1}{3})^2}\]

OpenStudy (anonymous):

we get \[\frac {\frac{2}{3}}{1-\frac{1}{9}}\]

OpenStudy (anonymous):

Yes, exactly! And I'm stuck. I got my answer as (18/27)/9 which I highly doubt is correct.

OpenStudy (anonymous):

\[=\frac{\frac{2}{3}}{\frac{8}{9}}\] \[=\frac{2}{3}\times \frac{9}{8}\] \[=\frac{3}{4}\] if my arithmetic is correct.

OpenStudy (anonymous):

How may I ask did you get 8/9?

OpenStudy (anonymous):

\[1-\frac{1}{9}=\frac{9}{9}-\frac{1}{9}=\frac{8}{9}\]

OpenStudy (anonymous):

lol aye yi yi, thank you so much. I really appreciate it.

OpenStudy (anonymous):

welcome!